OpenStudy (anonymous):

Find the number of units that produce a maximum revenue, R=95x-0.1x^2, where R is the total revenue in dollars and x is the number of units sold.

OpenStudy (anonymous):

To find the maximum revenue, we use calculus to take the derivative of the function and set the derivative equal to 0: R = 95x - 0.1x^2 R' = 95 - 0.2x 0 = 95 - 0.2x -95 = -0.2x x = 475 units Check my arithmetic, but that should be how you do it :)